OpenStudy (anonymous):

Write an equation of a hyperbola with vertices (3.-2) and (-9,-2) and foci (7,-2) and (-13,-2) a.(x-3)^2/36-(y-2)^2/64=1 b.(x-3)^2/12-(y-2)^2/16=1 c.(x+3)^2/12-(y+2)^2/16=1 d.(x+3)^2/36-(y+2)^2/64=1

OpenStudy (anonymous):

with the given vertices, can you find the center of the hyperbola?

OpenStudy (anonymous):

I dont know how to do this. Teacher was explaining it but I didnt understand

OpenStudy (anonymous):

the midpoint of the two vertices is the center of your hyperbola...

OpenStudy (anonymous):

at least find that so you can eliminate some choices...

OpenStudy (anonymous):

\[ \frac { (x-h)^2} {a^2}- \frac {(y-k)^2}{b^2} =1 \] (h,k) is the center. 2a = the distance between the two vertices 2 c=the distance between the two foci. \[ b^2 = c^2 -a^2 \]

OpenStudy (anonymous):

Write an equation of a hyperbola with vertices (3.-2) and (-9,-2) and foci (7,-2) and (-13,-2) center=( -3, -2)

OpenStudy (anonymous):

2a =12 a=6

OpenStudy (anonymous):

2c = 20 c=10

OpenStudy (anonymous):

\[ b^2 = 100 -36 =64 \]

OpenStudy (anonymous):

\[\frac {(x+3)^2}{36}-\frac {(y+2)^2}{64}=1 \]

OpenStudy (anonymous):

It is d.

OpenStudy (anonymous):

Thanks man.! Thats a bunch of steps thanks for being patient with me.

OpenStudy (anonymous):

yw